P. A copper penny has a mass of 2.5 g. Suppose it has a net charge of +1.2 nC. How many electrons has it lost? What fraction of the total number of electrons in an uncharged penny is this?
Answer #1

charge = 1.2 * 10^-9 C

charge on 1 electron = 1.6*10^-19 C

number of electron = (1.2*10^-9)/(1.6*10^-19)

= 7.5*10^9